# Expanding the Resolution Boundary of Outcome-Based Imperfect-Recall Abstraction in Games with Ordered Signals

### v1

@@ 6 Potential Outcome Isomorphism
Figure 3 illustrates the process of building POI in Leduc Hold’em. In the second stage, each hand corresponds to four possible opponent hole combinations, resulting in four signals per signal infoset. We tally the win-tie-loss outcomes for each signal in each infoset after showdown and group signal infosets with identical distributions into the same abstracted signal infoset, assigning a label to each. In the first stage, each hand will form different second-stage abstracted signal infosets as five different flop cards are dealt. We tally these distributions and consider hands with the same distribution as belonging to the same abstracted signal infoset. It can be observed that in Leduc Hold’em, the quantity of abstracted signal infosets in the first stage is equal to the quantity of signal infosets, indicating good identification capability. However, in the second stage, POI can only identify three abstracted signal infosets, significantly fewer than the nine signal infosets. In Leduc Hold’em, the quantity of signal infosets equals the quantity of lossless abstracted signal infosets. This also indicates that in Leduc Hold’em, POI has much lower resolution than LI. We will analyze this phenomenon in the next section.
@@ 7.2.1 E[HS] Algorithm
Here, $l,t,w$ represent the probabilities of signals in the infoset $\vartheta$ for player $i$ to lose, tie, and win in future showdown stages, respectively.
@@
Proof  We employ mathematical induction to prove that within the same potential outcome isomorphism class, different signal infosets will be assigned to the same abstracted signal infoset in the E[HS] algorithm for a set of parameters. To do this, it is sufficient to demonstrate that each signal infoset within the same potential outcome isomorphism class has the same equity value.
First, consider the last stage $r=|\mathfrak{r}|$. According to the expression in Equation (5), in this final stage of a 2-player scenario, the outcome isomorphism vector $(f_{i}^{(r),0}(\vartheta),f_{i}^{(r),1}(\vartheta),f_{i}^{(r),2}(\vartheta))$ for a signal infoset $\vartheta$ is simply the signal’s losing rate, tying rate, and winning rate for player $i$, each multiplied by the number of signals in $\vartheta$. This is in accordance with Equation (11), confirming our assumption holds.
Next, assuming the conclusion holds at stage $r+1$, let’s consider stage $r$. According to Algorithm 1, we can conclude that signal infosets within the same potential outcome isomorphism class $\tilde{\vartheta}^{(r)}$ share the same potential outcome feature. As stage $r$ is not the final stage, the potential outcome feature of a signal infoset $\vartheta$ at this stage can be represented as $(\tilde{\vartheta}^{(r+1)}_{1},\ldots,\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}})$, where $\tilde{\vartheta}^{(r+1)}_{1}\leq\ldots\leq\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}}$ are the indices of potential outcome isomorphism classes to which the signal infoset in $\tilde{\vartheta}^{(r)}$ can transition at stage $r+1$, sorted in ascending order. According to our assumption, at stage $r+1$, signal infosets within the same potential outcome isomorphism class have the same equity values. Therefore, for any potential outcome isomorphism class $\tilde{\vartheta}^{(r)}$ within stage $r$, the equity value vectors of signal infosets $\vartheta$ are the same and can be represented as
@@
Here, $e_{i}$ represents the equity value of $\tilde{\vartheta}^{(r+1)}_{i}$ in stage $r+1$. This completes the proof by mathematical induction, demonstrating that in all stages, the signal infosets within the same potential outcome isomorphism class have the same equity values.
@@ 7.2.2 PA&PAEMD Algorithm
The E[HS] algorithm, while proficient in reflecting the expected strength of hand rankings, lacks the ability to characterize the changes in hand strength across different stages of the game. In response to this limitation, Gilpin et al. (2007) introduced the PA algorithm. This algorithm employs a bottom-up approach, starting from the final stage, to perform k-means clustering on hands. Clustering in the final stage is based on the distances between cluster centroids and hand equities. For other stages, histograms are constructed for each cluster centroid to the next stage’s centroids, as well as for each hand to the next stage’s centroids. The clustering process utilizes the $L_{2}$ distance between the histograms of cluster centroids and hand histograms, ensuring a more nuanced understanding of hand strength dynamics throughout the game. Building upon the PA algorithm, Ganzfried and Sandholm (2014) further enhanced it with the PAEMD algorithm. This iteration incorporates the EMD between histograms of cluster centroids and hand histograms in non-final stages, alongside an efficient approximate method for calculating this distance.
@@
Proof  We first consider the last stage, $r=|\mathfrak{r}|$ It suffices to demonstrate that every signal infoset within the same potential outcome isomorphism class holds the same equity value, akin to the E[HS] algorithm’s case, i.e., the conclusion holds.
Next, assuming that the conclusion holds at stage $r+1$, it suffices at stage $r$ to prove that every distinct signal infoset within the same potential outcome isomorphism class has the same histogram of PA algorithm’s next-stage cluster centroids. For any potential outcome isomorphism class $\tilde{\vartheta}^{(r+1)}$ at stage $r$, any pair of signal infosets $\vartheta$ and $\vartheta^{\prime}$ from $\tilde{\vartheta}^{(r+1)}$ share an identical potential outcome feature. Since $r$ is not the final stage, this potential outcome feature is represented as a sparse histogram over potential outcome isomorphism classes from stage $r+1$, denoted as $(\tilde{\vartheta}^{(r+1)}_{1},\ldots,\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}})$. Consequently, there exist $\vartheta_{1},\vartheta_{1}^{\prime}\in\tilde{\vartheta}^{(r+1)}_{1},\ldots,\vartheta_{n^{(r+1)}},\vartheta_{n^{(r+1)}}^{\prime}\in\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}}$, where $\vartheta_{1},\ldots,\vartheta_{n^{(r+1)}}$ are signal infosets that $\vartheta$ can transition to in stage $r+1$, and $\vartheta_{1}^{\prime},\ldots,\vartheta_{n^{(r+1)}}^{\prime}$ are signal infosets that $\vartheta^{\prime}$ can transition to in stage $r+1$. According to the induction hypothesis, $\vartheta_{j}$ and $\vartheta_{j}^{\prime}$ are also assigned to the same cluster centroid $C_{j}$ under the PA algorithm with the given parameters $\psi$. Therefore, $\vartheta$ and $\vartheta^{\prime}$ share the same next-stage cluster centroid histogram in the PA algorithm with the given parameters $\psi$, and their sparse representations are both $(C_{1},\ldots,C_{n^{(r+1)}})$. We have now completed the proof by mathematical induction, demonstrating that across all stages, distinct signal infosets within the same potential outcome isomorphism class in the outcome isomorphism will be assigned to the same abstracted signal infoset in the PA algorithm with a set of parameters $\psi$.
@@ 7.3 Limitations of Potential Outcome Isomorphism
Through the above discussion, we have learned that POI is a common refinement for both the E[HS] algorithm and the PA&PAEMD algorithm. In other words, the strategies trained on the abstracted signal infosets generate by these two algorithms will not surpass those trained using POI, at least in one-player-abstraction-perspective. However, POI has its clear limitations. Table 1 displays the quantity of original signal infosets at each stage of Texas Hold’em, along with the quantity of abstracted signal infosets that both the LI and POI can identify at each stage. It can be observed that as the game progresses, the quantity of original signal infosets gradually increases, forming a triangular pattern. The LI algorithm can capture this pattern while simplifying the data. However, POI fails to reflect this pattern, and its quantity of abstracted signal infosets exhibits a spindle-shaped pattern, i.e., fewer in the early and late stages of the game and more in the middle stages. This indicates that POI excessively simplifies the data in the later stages of the game.
Stage Null Abstraction LI POI
@@
Let’s delve into the reasons behind this phenomenon. As illustrated in Figure 4, the factors influencing the scale of potential outcome isomorphism classes mainly revolve around two aspects. Firstly, the quantity of signal infosets to be identified grows as the game progresses. The greater the quantity of signal infosets, the more distinct potential outcome isomorphism classes can be identified. Secondly, the volume of information contained within the potential outcome features plays a pivotal role. Take the River stage as an example, the potential outcome features consider all possible showdown outcomes with opponent hands when drawing two private cards from the remaining 45 cards. At the Preflop stage, the potential outcome features encompass scenarios where, after removing one’s own two private cards from the remaining 52 cards, two private cards of the opponent and five community cards, resulting in all possible showdown outcomes. Due to POI being the type of imperfect-recall signal abstraction that demands players to forget all past information, the volume of information gradually diminishes as the game progresses. Consequently, the potential outcome features exhibit an increased probability of repetition due to reduced volume of information, thereby leading to a reduction in the scale of distinct potential outcome isomorphism classes identified. The combined influence of these two factors results in a trend where the quantity of potential outcome isomorphism classes follows a spindle-shaped distribution, and the phenomenon of excessive abstraction becomes more evident in the later stages.
@@ 8 K-Recall Outcome Isomorphism
2:$r\in\{1,\dots,|\mathfrak{r}|\}$. Stages.
@@
7:     Initialize $\mathcal{KC}_{i}^{(r)}$ vector as empty.
8:     Initialize $\mathcal{KD}_{i}^{(r)}$ array arbitrarily with length $|\varTheta_{i}^{(r)}|$.
@@
11:         Append $feature$ to $\mathcal{KC}_{i}^{(r)}$.
@@
13:     Eliminate duplicates from $\mathcal{KC}_{i}^{(r)}$.
14:     Sort the elements of $\mathcal{KC}_{i}^{(r)}$ in lexicographical order.
15:     Construct hash table $\mathcal{KCI}_{i}^{(r)}$ from $\mathcal{KC}_{i}^{(r)}$. Store the index $lexid$ and value $feature$ of $\mathcal{KC}_{i}^{(r)}$ in $\mathcal{KCI}_{i}^{(r)}$ as key-value pairs $(feature,lexid)$.
@@
19:         Update $\mathcal{KD}_{i}^{(r)}[idx]$ with $\mathcal{KCI}_{i}^{(r)}[feature]$.
@@
26:         $\vartheta_{j}\leftarrow$ the signal infoset predecessor in the $j$-th stage of $\vartheta$.
@@
We have identified the underlying reasons for the phenomenon of excessive abstraction in POI. This arises from the gradual reduction of information within potential outcome features, as discussed in the preceding section. During the identification of a signal infoset, only the information spanning from the current stage to the end of the game is employed, i.e., imperfect-recall that demands players to forget all past information. This, in turn, increases the probability of potential outcome feature repetition. In this section, we will introduce the concept of k-recall outcome isomorphism (KROI), which addresses this issue by supplementing information from past stages to mitigate the problem of excessive abstraction.
As illustrated in Algorithm 3, for a signal infoset at stage $r$, we can compute its $k$-recall outcome isomorphism, where $k=0,\dots,r-1$. The primary procedure of the algorithm closely mirrors that of the POI algorithm, with the only distinction being the replacement of the original potential outcome features with $k$-recall outcome features in lines 5 and 11. The $k$-recall outcome features for a signal infoset $\vartheta$ constitute a vector of length $k+1$, where the component at index $j$ stores the index of potential outcome isomorphism class for the predecessor signal infoset of $\vartheta$ at stage $r-j$. Since POI incorporates information from the current stage to the end of the game, $k$-recall outcome features encompass the information of $\vartheta$ spanning the preceding $k+1$ stages, including stage $r$.
@@
Figure 5 illustrates the process of constructing $1$-ROI in Leudc Hold’em. The bottom-up part follows the same procedure as constructing POI. In the first stage, only 00-ROI construction is possible, which aligns with the result of POI. The change occurs in the second stage, where each signal infoset is assigned the current stage’s POI class label, as well as the POI class label where its predecessor in the previous stage is located. These labels form a vector, known as the $1$-recall outcome feature. We group signal infosets with the same $1$-recall outcome feature into the same abstracted signal infoset. It can be observed that in Leduc Hold’em, the quantity of $1$-ROI classes in the second stage is 7, higher than the 3 abstracted signal infosets of POI. This represents an expansion of the resolution of the outcome-based imperfect-recall algorithm.
Table 2 presents the quantity of abstracted signal infosets identified by $k$-recall outcome isomorphism at each stage in HUNL&HULHE, varying with $k$. When $k=0$, $k$-recall outcome isomorphism is essentially equivalent to POI. For stages where $r\leq k$, the quantity of abstracted signal infosets is directly inherited from $(k-1)$-ROI. Furthermore, when incorporating information from all preceding stages (e.g., $3$-ROI in HUNL&HULHE), the quantity of abstracted signal infosets follows a triangular distribution, mirroring that observed in the LI and the original game’s signal infosets.
It should be noted that for games with a maximum stage of $r$, $r$-ROI serves as a perfect-recall signal abstraction. However, KROI can lead to more generalized outcome-based imperfect-recall signal abstraction algorithms, allowing players to retain partial memory instead of forgetting everything. For instance, when designing abstraction algorithms, we only cluster $\mathcal{D}_{i}^{(r)}$ at each stage $r$, without considering the connections between stages. Such an abstraction algorithm remains imperfect-recall, as it does not require consistent paths between abstracted signal infosets. Thus, we obtain a signal abstraction framework with higher resolution, with KROI serving as their resolution boundary.
Stage 00-ROI $1$-ROI $2$-ROI $3$-ROI

### v2

@@ 5 Potential Outcome Isomorphism
Figure 3 illustrates the process of building POI in Leduc Hold’em for player 1 (the result is similar for player 2). In the second phase, each hand corresponds to four possible opponent hole combinations, resulting in four signals per signal infoset. We tally the win-tie-loss outcomes for each signal in each infoset after showdown and group signal infosets with identical distributions into the same abstracted signal infoset, assigning a label to each. In the first phase, each hand will form different second-phase abstracted signal infosets as five different flop cards are dealt. We tally these distributions and consider hands with the same distribution as belonging to the same abstracted signal infoset. It can be observed that in Leduc Hold’em, the quantity of abstracted signal infosets in the first phase is equal to the quantity of signal infosets, indicating good identification capability. However, in the second phase, POI can only identify three abstracted signal infosets, significantly fewer than the nine signal infosets. In Leduc Hold’em, the quantity of signal infosets equals the quantity of lossless abstracted signal infosets. This also indicates that in Leduc Hold’em, POI has much lower resolution than LI. We will analyze this phenomenon in the next section.
@@ 6.2.1 E[HS] Algorithm
Here, $l,t,w$ represent the probabilities of signals in the infoset $\vartheta$ for player $i$ to lose, tie, and win in future showdown phases, respectively.
@@
Proof.
We employ mathematical induction to prove that within the same potential outcome isomorphism class, different signal infosets will be assigned to the same abstracted signal infoset in the E[HS] algorithm for a set of parameters. To do this, it is sufficient to demonstrate that each signal infoset within the same potential outcome isomorphism class has the same equity value.
First, consider the last phase $r=|\mathfrak{r}|$. According to the expression in Equation (5), in this final phase of a 2-player scenario, the outcome isomorphism vector $(f_{i}^{(r),0}(\vartheta),f_{i}^{(r),1}(\vartheta),f_{i}^{(r),2}(\vartheta))$ for a signal infoset $\vartheta$ is simply the signal’s losing rate, tying rate, and winning rate for player $i$, each multiplied by the number of signals in $\vartheta$. This is in accordance with Equation (11), confirming our assumption holds.
Next, assuming the conclusion holds at phase $r+1$, let’s consider phase $r$. According to Algorithm 1, we can conclude that signal infosets within the same potential outcome isomorphism class $\tilde{\vartheta}^{(r)}$ share the same potential outcome feature. As phase $r$ is not the final phase, the potential outcome feature of a signal infoset $\vartheta$ at this phase can be represented as $(\tilde{\vartheta}^{(r+1)}_{1},\ldots,\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}})$, where $\tilde{\vartheta}^{(r+1)}_{1}\leq\ldots\leq\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}}$ are the indices of potential outcome isomorphism classes to which the signal infoset in $\tilde{\vartheta}^{(r)}$ can transition at phase $r+1$, sorted in ascending order. According to our assumption, at phase $r+1$, signal infosets within the same potential outcome isomorphism class have the same equity values. Therefore, for any potential outcome isomorphism class $\tilde{\vartheta}^{(r)}$ within phase $r$, the equity value vectors of signal infosets $\vartheta$ are the same and can be represented as
@@
Here, $e_{i}$ represents the equity value of $\tilde{\vartheta}^{(r+1)}_{i}$ in phase $r+1$. This completes the proof by mathematical induction, demonstrating that in all phases, the signal infosets within the same potential outcome isomorphism class have the same equity values. ∎
@@ 6.2.2 PA&PAEMD Algorithm
The E[HS] algorithm, while proficient in reflecting the expected strength of hand rankings, lacks the ability to characterize the changes in hand strength across different phases of the game. In response to this limitation, Gilpin et al. [9] introduced the PA algorithm. This algorithm employs a bottom-up approach, starting from the final phase, to perform k-means clustering on hands. Clustering in the final phase is based on the distances between cluster centroids and hand equities. For other phases, histograms are constructed for each cluster centroid to the next phase’s centroids, as well as for each hand to the next phase’s centroids. The clustering process utilizes the $L_{2}$ distance between the histograms of cluster centroids and hand histograms, ensuring a more nuanced understanding of hand strength dynamics throughout the game. Building upon the PA algorithm, Ganzfried and Sandholm [12] further enhanced it with the PAEMD algorithm. This iteration incorporates the EMD between histograms of cluster centroids and hand histograms in non-final phases, alongside an efficient approximate method for calculating this distance.
@@
Proof.
We first consider the last phase, $r=|\mathfrak{r}|$ It suffices to demonstrate that every signal infoset within the same potential outcome isomorphism class holds the same equity value, akin to the E[HS] algorithm’s case in Section 6.2.1, i.e., the conclusion holds.
Next, assuming that the conclusion holds at phase $r+1$, it suffices at phase $r$ to prove that every distinct signal infoset within the same potential outcome isomorphism class has the same histogram of PA algorithm’s next-phase cluster centroids. For any potential outcome isomorphism class $\tilde{\vartheta}^{(r+1)}$ at phase $r$, any pair of signal infosets $\vartheta$ and $\vartheta^{\prime}$ from $\tilde{\vartheta}^{(r+1)}$ share an identical potential outcome feature. Since $r$ is not the final phase, this potential outcome feature is represented as a sparse histogram over potential outcome isomorphism classes from phase $r+1$, denoted as $(\tilde{\vartheta}^{(r+1)}_{1},\ldots,\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}})$. Consequently, there exist $\vartheta_{1},\vartheta_{1}^{\prime}\in\tilde{\vartheta}^{(r+1)}_{1},\ldots,\vartheta_{n^{(r+1)}},\vartheta_{n^{(r+1)}}^{\prime}\in\tilde{\vartheta}^{(r+1)}_{n^{(r+1)}}$, where $\vartheta_{1},\ldots,\vartheta_{n^{(r+1)}}$ are signal infosets that $\vartheta$ can transition to in phase $r+1$, and $\vartheta_{1}^{\prime},\ldots,\vartheta_{n^{(r+1)}}^{\prime}$ are signal infosets that $\vartheta^{\prime}$ can transition to in phase $r+1$. According to the induction hypothesis, $\vartheta_{j}$ and $\vartheta_{j}^{\prime}$ are also assigned to the same cluster centroid $C_{j}$ under the PA algorithm with the given parameters $\psi$. Therefore, $\vartheta$ and $\vartheta^{\prime}$ share the same next-phase cluster centroid histogram in the PA algorithm with the given parameters $\psi$, and their sparse representations are both $(C_{1},\ldots,C_{n^{(r+1)}})$. We have now completed the proof by mathematical induction, demonstrating that across all phases, distinct signal infosets within the same potential outcome isomorphism class in the outcome isomorphism will be assigned to the same abstracted signal infoset in the PA algorithm with a set of parameters $\psi$. ∎
@@ 6.3 Limitations of Potential Outcome Isomorphism
Through the above discussion, we have learned that POI is a common refinement for both the E[HS] algorithm and the PA&PAEMD algorithm. In other words, the strategies trained on the abstracted signal infosets generate by these two algorithms will not surpass those trained using POI, at least in one-player-abstraction-perspective. However, POI has its clear limitations. Table 1 displays the quantity of original signal infosets at each phase of Texas Hold’em, along with the quantity of abstracted signal infosets that both the LI and POI can identify at each phase. It can be observed that as the game progresses, the quantity of original signal infosets gradually increases, forming a triangular pattern. The LI algorithm can capture this pattern while simplifying the data. However, POI fails to reflect this pattern, and its quantity of abstracted signal infosets exhibits a spindle-shaped pattern, i.e., fewer in the early and late phases of the game and more in the middle phases. This indicates that POI excessively simplifies the data in the later phases of the game.
Phase Null Abstraction LI POI
@@
Let’s delve into the reasons behind this phenomenon. As illustrated in Figure 4, the factors influencing the scale of potential outcome isomorphism classes mainly revolve around two aspects. Firstly, the quantity of signal infosets to be identified grows as the game progresses. The greater the quantity of signal infosets, the more distinct potential outcome isomorphism classes can be identified. Secondly, the volume of information contained within the potential outcome features plays a pivotal role. Take the River phase as an example, the potential outcome features consider all possible showdown outcomes with opponent hands when drawing two private cards from the remaining 45 cards. At the Preflop phase, the potential outcome features encompass scenarios where, after removing one’s own two private cards from the remaining 52 cards, two private cards of the opponent and five community cards, resulting in all possible showdown outcomes. Due to POI being the type of imperfect-recall signal abstraction that demands players to forget all past information, the volume of information gradually diminishes as the game progresses. Consequently, the potential outcome features exhibit an increased probability of repetition due to reduced volume of information, thereby leading to a reduction in the scale of distinct potential outcome isomorphism classes identified. The combined influence of these two factors results in a trend where the quantity of potential outcome isomorphism classes follows a spindle-shaped distribution, and the phenomenon of excessive abstraction becomes more evident in the later phases.
@@ 7 K-Recall Outcome Isomorphism
2:$r\in\{1,\dots,|\mathfrak{r}|\}$. Phases.
@@
7:     Initialize $\mathcal{KC}_{i}^{(r,k)}$ vector as empty.
8:     Initialize $\mathcal{KD}_{i}^{(r,k)}$ array arbitrarily with length $|\varTheta_{i}^{(r)}|$.
@@
11:         Append $feature$ to $\mathcal{KC}_{i}^{(r,k)}$.
@@
13:     Eliminate duplicates from $\mathcal{KC}_{i}^{(r,k)}$.
14:     Sort the elements of $\mathcal{KC}_{i}^{(r,k)}$ in lexicographical order.
15:     Construct hash table $\mathcal{KCI}_{i}^{(r,k)}$ from $\mathcal{KC}_{i}^{(r,k)}$. Store the index $lexid$ and value $feature$ of $\mathcal{KC}_{i}^{(r,k)}$ in $\mathcal{KCI}_{i}^{(r,k)}$ as key-value pairs $(feature,lexid)$.
@@
19:         Update $\mathcal{KD}_{i}^{(r,k)}[idx]$ with $\mathcal{KCI}_{i}^{(r,k)}[feature]$.
@@
26:         $\vartheta_{j}\leftarrow$ the signal infoset predecessor in the $j$-th phase of $\vartheta$.
@@
We have identified the underlying reasons for the phenomenon of excessive abstraction in POI. This arises from the gradual reduction of information within potential outcome features, as discussed in the preceding section. During the identification of a signal infoset, only the information spanning from the current phase to the end of the game is employed, i.e., imperfect-recall that demands players to forget all past information. This, in turn, increases the probability of potential outcome feature repetition. In this section, we will introduce the concept of k-recall outcome isomorphism (KROI), which addresses this issue by supplementing information from past phases to mitigate the problem of excessive abstraction.
As illustrated in Algorithm 3, for a signal infoset at phase $r$, we can compute its $k$-recall outcome isomorphism, where $k=0,\dots,r-1$. The primary procedure of the algorithm closely mirrors that of the POI algorithm, with the only distinction being the replacement of the original potential outcome features with $k$-recall outcome features in lines 5 and 11. The $k$-recall outcome features for a signal infoset $\vartheta$ constitute a vector of length $k+1$, where the component at index $j$ stores the index of potential outcome isomorphism class for the predecessor signal infoset of $\vartheta$ at phase $r-j$. Since POI incorporates information from the current phase to the end of the game, $k$-recall outcome features encompass the information of $\vartheta$ spanning the preceding $k+1$ phases, including phase $r$.
@@
Figure 5 illustrates the process of constructing $1$-ROI in Leudc Hold’em. The bottom-up part follows the same procedure as constructing POI. In the first phase, only 00-ROI construction is possible, which aligns with the result of POI. The change occurs in the second phase, where each signal infoset is assigned the current phase’s POI class label, as well as the POI class label where its predecessor in the previous phase is located. These labels form a vector, known as the $1$-recall outcome feature. We group signal infosets with the same $1$-recall outcome feature into the same abstracted signal infoset. It can be observed that in Leduc Hold’em, the quantity of $1$-ROI classes in the second phase is 7, higher than the 3 abstracted signal infosets of POI. This represents an expansion of the resolution of the outcome-based imperfect-recall algorithm.
Table 2 presents the quantity of abstracted signal infosets identified by $k$-recall outcome isomorphism at each phase in HUNL&HULHE, varying with $k$. When $k=0$, $k$-recall outcome isomorphism is essentially equivalent to POI. For phases where $r\leq k$, the quantity of abstracted signal infosets is directly inherited from $(k-1)$-ROI. Furthermore, when incorporating information from all preceding phases (e.g., $3$-ROI in HUNL&HULHE), the quantity of abstracted signal infosets follows a triangular distribution, mirroring that observed in the LI and the original game’s signal infosets.
It should be noted that for games with a maximum phase of $r$, $r$-ROI serves as a perfect-recall signal abstraction. However, KROI can lead to more generalized outcome-based imperfect-recall signal abstraction algorithms, allowing players to retain partial memory instead of forgetting everything. For instance, when designing abstraction algorithms, we only cluster $\mathcal{D}_{i}^{(r)}$ at each phase $r$, without considering the connections between phases. Such an abstraction algorithm remains imperfect-recall, as it does not require consistent paths between abstracted signal infosets. Thus, we obtain a signal abstraction framework with higher resolution, with KROI serving as their resolution boundary.
Phase 00-ROI $1$-ROI $2$-ROI $3$-ROI